Water with an alkalinity of 4.26 meq/L, is to be soften by the lime-soda process has the following components: 3.25 meq/L calcium, 2.67 meq/L magnesium, and 0.36 meq/L carbon dioxide. The amount of calcium oxide (85% pure), in lb/MG, needed per million gallons is
Answer #1

1 gallon = 3.785 liters

Million Gallons = 106 gallons = 3.785*106 liters => 1 liter = 1/(3.785*106) million gallons

1 Kg = 2.20462 lb => 1 mg = 10-6kg = 2.20462* 10-6 lb

Lime required = 281.1 mg/l = 281.1*2.20462* 10-6*3.785*106 = 2345.635 lb/MG (ans)

I hope its clear!

Please rate.

Thank you!